Given the equation f=TIy, where X-27.11±0.04 andy=0.0899±0.0005 What is the absolute uncertainty in f? Select one: a 7.66±0.05 b.7.6610.08 с 7.657 0.007 07,66士0.04
Answer #1

Here (pi)xy = 27.11×0.0899=7.66

Now error = 7.66×(0.04/27.11 +0.0005/0.0899)

= 0.054

Hence f= 7.66 ± 0.05

Hence first is correct